OpenStudy (anonymous):

What is the relation between 220 & 284 ?

OpenStudy (zehanz):

There are all kinds of relations: They are both positive They are both integer They are both even Could you be more specific, or are these the kind of relations you are looking for?

OpenStudy (anonymous):

You are right , but the relation between 220 and 284 is "FRIEND" . That means 220 & 284 is called "Friend" according to Pythgoras opinion.

OpenStudy (shubhamsrg):

-_-

OpenStudy (zehanz):

This is what it means: (sum of dividers of 220) = 1 + 2 + 4 + 5 + 10 + 11 + 20 + 22 + 44 + 55 + 110 = 284 ((sum of dividers of 284) = 1 + 2 + 4 + 71 + 142 = 220

OpenStudy (shubhamsrg):

ohh..thats cool.

OpenStudy (anonymous):

These numbers are amicable pairs.

OpenStudy (zehanz):

amicable, friends, why not lovers!